# Reassessing the gallium anomaly using exact electron wave functions

### v2

@@ Abstract
The gallium anomaly, a persistent discrepancy exceeding $4\sigma$ in the 71Ga neutrino capture rates from 51Cr and 37Ar calibration sources by the GALLEX, SAGE, and recently BEST experiments, has challenged particle physics and nuclear theory for over three decades. We present a new calculation for the neutrino capture cross-section, abandoning the conventional leading-order approximation for electronic wave functions by numerically solving the Dirac-Coulomb equation for both bound and continuum electron states. Finally, we re-evaluate the gallium anomaly, updating its global significance and presenting the most up-to-date status of its interpretation in terms of sterile neutrinos.
@@ II Determination of the neutrino capture cross section
In the literature, it is common practice to assume that the neutrino can be considered to be a plane wave normalized in such a way to cancel $\Omega_{\nu}$ in Eq. (1), and that the electron wave functions vary slowly within the nuclear volume. Therefore, the latter can be factored out by considering its value at a specific distance $r_{0}$, usually chosen to be $r_{0}=0$ or $r_{0}=R_{\rm box}$, where $R_{\rm box}$ represents the equivalent box density radius, namely $R_{\rm box}=\sqrt{5/3}\,R_{\rm ch}$, with $R_{\rm ch}$ the nuclear charge radius. In this way, one can isolate the IBD nuclear matrix element, $\mathcal{M}^{\rm IBD}_{\rm nuc}$, [^3] In the literature, the nuclear matrix element is often referred to in terms of the Gamow-Teller strength $B_{\rm GT}=|\mathcal{M}_{\rm nuc}|^{2}/(2J+1)$.
@@
with $J_{\text{Ge}}=1/2$ being the nuclear spin of germanium, $n_{x}$ the relative occupation number of electrons in the atomic $x$ shell, and $\mathcal{H}^{\rm EC}_{x}$ the EC matrix element. In both the IBD cross section in Eq. (1) and the EC rate in Eq. (4) we account for radiative corrections and the weak magnetism contribution following Ref. [27]. Factorizing out the leptonic wave functions as in the IBD case, the electron-capture matrix element becomes
@@
with $\epsilon_{\text{o}}^{x}$ accounting for exchange and overlap effects, and $P_{\text{L,M,K}}$ being the experimentally measured electron-capture probabilities for the L, M and K shells [41, 46, 37]. Here, $|\psi^{\rm b}_{e,1s}(r_{0})|^{2}$ is the $1s$ electron density at the nucleus obtained from
@@
Following the detailed balance principle, $|\mathcal{M}^{\rm IBD}_{\rm nuc}|^{2}\stackrel{{\scriptstyle\rm db}}{{=}}|\mathcal{M}^{\rm EC}_{\rm nuc}|^{2}$, the ground-state neutrino-induced IBD cross section reduces to [27]
@@
where, to match the usual conventions, we have normalized the non-distorted electron wave function, $|\psi_{e}(r_{0})|_{\text{free}}^{2}$, to cancel $\Omega_{e}$ in Eq. (1), and introduced the Fermi function, $\mathcal{F}(E_{e},Z)=|\psi_{e}(r_{0})|^{2}/|\psi_{e}(r_{0})|^{2}_{\text{free}}$, which corrects for the distortion of the electron wave function due to the Coulomb potential of the daughter nucleus with $Z$ being its atomic number. Here, we use a generalised Fermi function given by [54]
@@
In this work, we developed a software based on the RADIAL package [53] that allows us to numerically evaluate the Fermi function in Eq. (10), which requires the use of the exact Dirac-Coulomb radial electron wave functions, $\psi_{e,\kappa}(\mathbf{r})$, that are given in terms of the large and small components, $g_{\kappa}(r)$ and $f_{\kappa}(r)$, obtained by resolving
@@
where $V(r)$ is the atomic potential of the germanium atom. Similar functions, $\psi^{\rm b}_{e,n\kappa}(\mathbf{r})$, $g_{n\kappa}(r)$ and $f_{n\kappa}(r)$, can be defined for the case of bound electron states, with the replacement of the electron energy $E_{e}$ with $E_{e}^{nk}=m_{e}-|E_{\rm bind}^{nk}|$, where $E_{\rm bind}^{nk}$ is the binding energy and introducing a subscript $n$ indicating the principal quantum number. In our work, we consider the DHFS potential, which is given by
@@
where $V_{\text{nuc}}(r)$ is the nuclear potential. The latter has been determined by considering a two-parameter Fermi (2pF) distribution [55] with $R_{\rm ch}=4.032(2)$ fm [44] and skin thickness $t=2.3\,\mathrm{fm}$. The electronic potential $V_{\text{el}}(r)$ accounts for the interaction energy of an electron with the entire electronic cloud at a distance $r$, and the exchange potential $V_{\text{ex}}(r)$ is defined as
@@
where the constant $C_{\text{ex}}$ assumes different values according to different calculations, namely $C_{\text{ex}}=3/2$ for Slater [56, 53] and $C_{\text{ex}}=1$ for Kohn-Sham [57, 53]. [^4] The choice of this coefficient results in differences up to $1\%$ on the electron density at the nucleus [58]. In our calculation, we considered $C_{\text{ex}}=3/2$. To ensure the correct long-range behavior, we apply the so-called Latter tail correction [59, 53]. The electron density, $\rho(r)$, which enters both $V_{\rm el}(r)$ and $V_{\rm ex}(r)$, is determined through a self-consistent solution of the DHFS equations iteratively [53].
The generalised definition in Eq. (10) has to be compared to the common definition of the Fermi function which can be found in the literature (see e.g. Ref. [27]), obtained by the product of various contributions, namely
@@
$F_{0}(E_{e},Z)=4(2p_{e}R_{\rm box})^{2\gamma-2}\frac{\absolutevalue{\Gamma(\gamma+i\eta)}^{2}}{\absolutevalue{\Gamma(2\gamma+1)}^{2}}\text{e}^{\pi\eta}\,,$ (15)
with $\gamma\equiv\sqrt{1-(\alpha Z)^{2}}$ and $\eta\equiv(\alpha ZE_{e})/p_{e}$, where $\alpha$ is the fine-structure constant. The additional term $L_{0}(E_{e},Z)$ corrects for the finite size of the nucleus considering a box density with equivalent radius $R_{\rm box}=\sqrt{5/3}R_{\rm ch}$, while $U(E_{e},Z)$ is introduced to improve the nuclear distribution from a box density to a more realistic Fermi distribution. Finally, $S(E_{e},Z)$ introduces a correction to the outgoing wave function due to atomic screening [54, 60, 53]. We have checked that by considering a nucleus with a uniform density and evaluating the Fermi function for $r=0$, we retrieve the tabulated values of the Fermi function from Behrens et al. [38], expressed as the first two terms of Eq. (14), $F_{0}L_{0}$.
@@
In Fig. 1, we show our result for the Fermi function calculated using the exact DHFS $g_{\kappa}(r)$ and $f_{\kappa}(r)$ wave functions in Eq. (10). Our result has been obtained by calculating all of the contributions to the Fermi function within a unique and coherent framework and produce slightly higher values compared to those reported in Ref. [27], as shown in Table 1 for $r=0$.
Besides the numerical evaluation of the Fermi function, we employed the same dedicated software, based on the RADIAL package [53], also to numerically evaluate the bound electron wave functions in Eq. (8) for a finite-size nucleus. For this, we use a 2pF distribution with the first measurement of the charge radius of ${}^{71}{\textrm{Ge}}$ [44] as an input parameter, equal to $R_{\rm ch}=4.032(2)$ fm. In Fig. 1, we show the electron density for the $1s$ bound state of 71Ge as obtained by our numerical calculation. In Table 2, we compare the electron densities obtained in this work with those reported in the literature for different methods and assumptions. For a better comparison, we also report the values we would obtain with our software using the same underlying assumptions. The electron densities evaluated with our code are compatible with those reported in the literature within about 3%, when using the same inputs.
@@
As a further improvement, we move beyond the conventional approximation of evaluating the electron wave function at a single point, as done in Eqs. (2) and (5). We propose the adoption of averaged wave functions over the ${}^{71}{\textrm{Ge}}$ nuclear charge density, in both the calculations of IBD and EC amplitudes. [^5] The average should be evaluated over the Gamow-Teller transition density, defined as the overlap between the 71Ge and 71Ga nuclear wave functions with the interaction Hamiltonian. However, this quantity is not currently accessible from first principles in nuclear theory, and we leave its inclusion for an upcoming work. This modification also results in an averaging procedure of the Fermi function in Eq. (10), ¯$\overline{\mathcal{F}}(E^{j}_{e},Z_{\text{Ge}})$, consistent with the prescription originally suggested by Bahcall [35] for a point-like nucleus. [^6] Bahcall suggested to apply the averaging procedure to $F_{0}$ considering that the “electron-capture can occur anywhere within the nuclear volume” [35]. This is not completely justified, since in the evaluation of $F_{0}$ the nucleus is assumed to be point-like and considering that the Fermi function corrects the wave function of the outgoing electron and it does not refer to the electron captured in the EC process. On the contrary, the introduction of exact electron wave functions, as done in this work, naturally leads to the evaluation of the average of the latter within the nucleus when factoring out the electron wave function from the matrix element as in Eq. (2). The averaging is done through
¯$\overline{\mathcal{F}}(E_{e},Z)=\frac{\int{\rho_{\rm ch}(r)\,\mathcal{F}}(E_{e},Z,r)\,\mathrm{d}^{3}r}{\int{\rho_{\rm ch}(r)\,\mathrm{d}^{3}r}}\,,$ (16)
@@
The effect of averaging the electron density is shown in the last three rows of Table 2, where we compare the values obtained at a fixed radius, namely $r_{0}=0$ and $r_{0}=\sqrt{5/3}R_{\rm ch}$, with that retrieved after the averaging procedure.
@@
51Cr 1.031 2.774 2.806 2.736
1.030 2.775 2.806 2.738
1.025 2.779 2.810 2.742
0.711 3.306 3.359 3.281
0.710 3.309 3.361 3.284
0.705 3.330 3.384 3.305
37Ar 1.092 2.734 2.763 2.696
1.092 2.734 2.764 2.696
1.089 2.736 2.765 2.698
Table 1: Comparison between the Fermi function defined according to Eq. (14[27], $F(E_{e},Z_{\text{Ge}})$, and those calculated with our software according to Eqs. (10) and (16), $\mathcal{F}(E_{e},Z_{\text{Ge}})$ and ¯$\overline{\mathcal{F}}(E_{e},Z_{\text{Ge}})$, for electron energies $E_{e}$ relevant to ${}^{51}\text{Cr}$ and ${}^{37}\text{Ar}$ neutrino sources. In the first two cases we use $r_{0}=0$.
@@
To calculate $f_{\text{EC}}$, following Eq. (7), we keep into account the so-called overlap and exchange correction, $\epsilon_{o}^{1s}$, given that we write the total rate in terms of the 1s-capture rate. We compute the average of the values reported in Refs. [63, 64], as previously done in Ref. [27], yielding to $\epsilon_{o}^{1s}=-0.013\pm 0.014$. We also consider an additional term due to the weak magnetism correction, $[1+\epsilon_{\text{q}}]_{\text{EC}}$, with $\epsilon_{\text{q}}=(4.9\pm 1.5)\,\times 10^{-4}$ [27].
The values obtained in this work for $f_{\rm EC}$ are reported in Table 2 for different calculations of the bound electron wave functions.
[…]
Band et al. (1979) [65] ($r_{0}=0$) Box [A=74] $1.2\,\text{A}^{1/3}$ 0.0692 0.0678 0.0226
Band et al. (1986) [58] ($r_{0}=0$) for $C_{\rm ex}=\frac{3}{2}$ [^7] Here, the values of $C_{\rm ex}$ refer to the potential in Eq. (13). Box [A=74] $1.201\,\text{A}^{1/3}$ 0.0696 0.0677 0.0225
Band et al. (1986) [58] ($r_{0}=0$) for DF [^8] Here, DF refers to the use of a different resolution method, namely Dirac-Fock instead of the DHFS one, for which the exchange effect is evaluated without any approximation. Box [A=74] $1.201\,\text{A}^{1/3}$ 0.0692
Behrens et al. [38] ($r_{0}=0$) Box [A=69] $1.2\,\text{A}^{1/3}$ 0.0694 0.0678 0.0226
Bambynek el al. [37] ($r_{0}=0$) 2pF [A=71] $c=4.43,\,t=2.4$ 0.0692 0.0677 0.0225
Band et al. (1979) [65] ($r_{0}=5.04\,\mathrm{fm}$) Box [A=74] $1.2\,\text{A}^{1/3}$ 0.0675 0.0661 0.0220
Behrens et al. [38] ($r_{0}=4.92\,\mathrm{fm}$) Box [A=69] $1.2\,\text{A}^{1/3}$ 0.0677 0.0662 0.0220
Martin et al. [66] ($r_{0}=4.49\,\mathrm{fm}$) [^9] Ref.[66] quotes $|g_{1s}|^{2}$ and does not report the value of A. Thus, we consider A=74 which corresponds to the most abundant stable isotope of germanium. 2pF [A=74] $c=4.49,\,t=2.5$ 0.0676 0.0663 0.0221
Elliott et al. [27] (avg.) [^10] Elliott et al. use the same wave functions used by Bahcall in Ref. [26], after applying, in their calculations, corrections to the neutrino energy. However, there is not enough information to compare our calculation to their result. 2pF [A=71] 0.0679
Cadeddu et al. ($r_{0}=0$) 2pF [A=71] $c=4.559,\,t=2.3$ 0.0678 0.0226
Cadeddu et al. ($r_{0}=5.20\,\text{fm}$) 2pF [A=71] $c=4.559,\,t=2.3$ 0.0660 0.0219
Cadeddu et al. (avg.) 2pF [A=71] $c=4.559,\,t=2.3$ 0.0666 0.0222
Table 2: Comparison of different electron densities for the $1s$ bound state of 71Ge at a specific radius $r_{0}$, $4\pi\,|\psi^{\rm b}_{e,1s}(r_{0})|^{2}$, calculated according to various nuclear charge densities, $\rho_{\rm ch}(r)$. “A” refers to the considered germanium mass number which typically enters in the definition of the charge radius. The table presents results from various references, specifying $\rho_{\rm ch}(r)$ (box or 2pF distribution), the associated input parameters, and the corresponding values of $4\pi\,|\psi^{\rm b}_{e,1s}(r_{0})|^{2}$ both from the reference and from our recalculation. In the last three rows, we show the values obtained in this work for $r_{0}=0$, $r_{0}=\sqrt{5/3}R_{\rm ch}$ and with the averaging procedure (avg). In the last column we report the corresponding values of $f_{\text{EC}}$ calculated within our framework.
@@
$B\mathrm{{}_{GT}}(\mathrm{gs})=\frac{|\mathcal{M}_{\rm nuc}^{\rm{IBD}}|^{2}}{2J_{\rm Ga}+1}=0.0878\pm 0.0012\,,$ (23)
@@
In Table 3, we report the ground-state cross section for the IBD process for 51Cr and 37Ar neutrino sources. The average procedure reduces the ground state cross section by 3.3%, indicating a non-negligible correction that should be taken into account.

### v3

# Reassessing the gallium anomaly using self-consistent electron wave functions

@@ Abstract
The gallium anomaly, a persistent discrepancy exceeding $4\sigma$ in the 71Ga neutrino capture rates from 51Cr and 37Ar radioactive sources by the GALLEX, SAGE, and recently BEST experiments, has challenged particle physics and nuclear theory for over three decades. We present a new calculation of the neutrino capture cross-section, abandoning the conventional leading-order approximation for electronic wave functions by numerically solving the Dirac-Coulomb equation for both bound and continuum electron states. Finally, we re-evaluate the gallium anomaly, updating its global significance and presenting the most up-to-date status of its interpretation in terms of sterile neutrinos.
@@ II Determination of the neutrino capture cross section
In the literature, it is common practice to assume that the neutrino can be considered to be a plane wave normalized in such a way to cancel $\Omega_{\nu}$ in Eq. (1), and that the electron wave functions vary slowly within the nuclear volume. Therefore, the latter can be factored out by considering its value at a specific distance $r_{0}$, usually chosen to be $r_{0}=0$ or $r_{0}=R_{\rm box}$, where $R_{\rm box}$ represents the equivalent box density radius, namely $R_{\rm box}=\sqrt{5/3}\,R_{\rm ch}$, with $R_{\rm ch}$ the nuclear charge radius. In this way, one can isolate the IBD nuclear matrix element, $\mathcal{M}^{\rm IBD}_{\rm nuc}$, [^3] In the literature, the nuclear matrix element is often referred to in terms of the Gamow-Teller strength $B_{\rm GT}=|\mathcal{M}_{\rm nuc}|^{2}/(2J+1)$, $J$ being the total angular momentum of the initial nuclear state.
@@
with $J_{\text{Ge}}=1/2$ the nuclear spin of the ground state of $^{71}{\textrm{Ge}}$, $n_{x}$ the relative occupation number of electrons in the atomic $x$ shell, and $\mathcal{H}^{\rm EC}_{x}$ the EC matrix element. In both the IBD cross section in Eq. (1) and the EC rate in Eq. (4) we account for radiative corrections and the weak magnetism contribution following Ref. [27]. Factorizing out the leptonic wave functions as in the IBD case, the electron-capture matrix element becomes
@@
with $\epsilon_{\text{o}}^{1s}$ accounting for exchange and overlap effects, and $P_{\text{L,M,K}}$ being the experimentally measured electron-capture probabilities for the L, M and K shells [41, 46, 37]. Here, $|\psi^{\rm b}_{e,1s}(r_{0})|^{2}$ is the $1s$ electron density at the nucleus obtained from
@@
Following the detailed balance principle, $|\mathcal{M}^{\rm IBD}_{\rm nuc}|^{2}\mathrel{{\mathop{=}\limits^{\rm db}}}|\mathcal{M}^{\rm EC}_{\rm nuc}|^{2}$, the ground-state neutrino-induced IBD cross section reduces to [27]
@@
where, to match the usual conventions, we have normalized the non-distorted electron wave function, $|\psi_{e}(r_{0})|_{\text{free}}^{2}$, to cancel $\Omega_{e}$ in Eq. (1), and introduced the Fermi function, $\mathcal{F}(E_{e},Z)=|\psi_{e}(r_{0})|^{2}/|\psi_{e}(r_{0})|^{2}_{\text{free}}$, which corrects for the distortion of the electron wave function due to the Coulomb potential of the daughter nucleus, with $Z$ being its atomic number. Here, we use a generalised Fermi function given by [54]
@@
In this work, we developed a software based on the RADIAL package [53] that allows us to numerically evaluate the Fermi function in Eq. (10), using the exact DHFS radial electron wave functions, $\psi_{e,\kappa}(\mathbf{r})$, that are given in terms of the large and small components, $g_{\kappa}(r)$ and $f_{\kappa}(r)$, obtained by solving
@@
where $V(r)$ is the atomic potential of the germanium atom.
Figure 1: Electron density at the nucleus as a function of the radial distance from the origin. We compare our calculations based on RADIAL [53] (labelled “Cadeddu et al.”) with that obtained using GRASP [55] and with those available in the literature [56, 57, 58, 38]. The horizontal dashed lines indicate the values obtained after applying the averaging procedure. The vertical dotted and dash-dotted lines represent $R_{\mathrm{ch}}$ and $R_{\mathrm{box}}=\sqrt{5/3}\,R_{\rm ch}$, respectively.
Similar functions, $\psi^{\rm b}_{e,n\kappa}(\mathbf{r})$, $g_{n\kappa}(r)$ and $f_{n\kappa}(r)$, can be defined for the case of bound electron states, with the replacement of the electron energy $E_{e}$ with $E_{e}^{n\kappa}=m_{e}-|E_{\rm bind}^{n\kappa}|$, where $E_{\rm bind}^{n\kappa}$ is the binding energy and introducing a subscript $n$ indicating the principal quantum number. In our work, we consider the DHFS potential, which is given by
@@
to which we apply the so-called Latter tail correction [59, 53], in order to ensure the correct long-range behavior. Here, $V_{\text{nuc}}(r)$ represents the nuclear potential, which has been determined by considering a two-parameter Fermi (2pF) distribution [60]with the first measurement of the charge radius of $^{71}{\textrm{Ge}}$ as an input parameter, equal to $R_{\rm ch}=4.032\pm 0.002$ fm [44], and skin thickness $t=2.3\,\mathrm{fm}$. The electronic potential $V_{\text{el}}(r)$ accounts for the interaction energy of an electron with the entire electronic cloud at a distance $r$, and the exchange potential $V_{\text{ex}}(r)$ is defined as
@@
where $\rho(r)$, which enters both $V_{\rm el}(r)$ and $V_{\rm ex}(r)$, is determined through a self-consistent solution of the DHFS equations iteratively [53]. The constant $C_{\text{ex}}$, as discussed in Ref. [61, 62], varies according to the chosen potential model. For example, $C_{\text{ex}}=3/2$ corresponds to the Slater approximation [63, 53] while $C_{\text{ex}}=1$ to the Kohn-Sham one [64, 53].
Reference $\rho_{\rm ch}(r)$ $\rho_{\rm ch}(r)$ input par.[fm]\,[$\text${fm}] $4\pi\,|\psi^{\rm b}_{e,1s}(r_{0})|^{2}$
Band et al. (1979) [58] ($r_{0}=0$) DF [^4] Here, DF refers to the use of a different resolution method, namely Dirac-Fock instead of the DHFS one, for which the exchange effect is evaluated without any approximation. The same result has also been reported in Ref. [57]. Box [A=74] $1.2\,\text{A}^{1/3}$ 0.0692
Band et al. (1986) [57] ($r_{0}=0$) for $C_{\rm ex}=\frac{3}{2}$ Box [A=74] $1.201\,\text{A}^{1/3}$ 0.0696
Behrens et al. [38] ($r_{0}=0$) Box [A=69] $1.2\,\text{A}^{1/3}$ 0.0694
Bambynek el al. [37] ($r_{0}=0$) 2pF [A=71] c=4.43,t=2.4c=4.43,\,t=2.4 0.0692
Band et al. (1979) [58] ($r_{0}=5.04\,\mathrm{fm}$) DF Box [A=74] $1.2\,\text{A}^{1/3}$ 0.0675
Behrens et al. [38] ($r_{0}=4.92\,\mathrm{fm}$) Box [A=69] $1.2\,\text{A}^{1/3}$ 0.0677
Martin et al. [56] ($r_{0}=4.49\,\mathrm{fm}$) [^5] Ref. [56] quotes $|g_{1s}|^{2}$ and does not report the value of A. Thus, we consider A=74 which corresponds to the most abundant stable isotope of germanium. 2pF [A=74] c=4.49,t=2.5c=4.49,\,t=2.5 0.0676
Elliott et al. [27] (avg.) [^6] Ref. [27] uses the same wave functions reported by Bahcall in Ref. [26], after applying, in their calculations, corrections to the neutrino energy. 2pF [A=71] 0.0679
GRASP [55] ($r_{0}=0$) 2pF [A=71] c=4.559,t=2.3c=4.559,\,t=2.3 0.0689
GRASP [55] ($r_{0}=5.20\,\text{fm}$) 2pF [A=71] c=4.559,t=2.3c=4.559,\,t=2.3 0.0671
GRASP [55] (avg.) 2pF [A=71] c=4.559,t=2.3c=4.559,\,t=2.3 0.0678
Cadeddu et al. ($r_{0}=0$) for $C_{\rm ex}=\frac{3}{2}$ 2pF [A=71] c=4.559,t=2.3c=4.559,\,t=2.3 0.0695
Cadeddu et al. ($r_{0}=0$) for $C_{\rm ex}=1$ 2pF [A=71] c=4.559,t=2.3c=4.559,\,t=2.3 0.0689
Cadeddu et al. ($r_{0}=0$) for $C_{\rm ex}=1.25$ 2pF [A=71] c=4.559,t=2.3c=4.559,\,t=2.3 0.0692
Cadeddu et al. ($r_{0}=5.20\,\text{fm}$) for $C_{\rm ex}=\frac{3}{2}$ 2pF [A=71] c=4.559,t=2.3c=4.559,\,t=2.3 0.0677
Cadeddu et al. ($r_{0}=5.20\,\text{fm}$) for $C_{\rm ex}=1$ 2pF [A=71] c=4.559,t=2.3c=4.559,\,t=2.3 0.0671
Cadeddu et al. ($r_{0}=5.20\,\text{fm}$) for $C_{\rm ex}=1$.25 2pF [A=71] c=4.559,t=2.3c=4.559,\,t=2.3 0.0674
Cadeddu et al. (avg.) for $\mathbf{C_{\rm ex}=1.25}$ 2pF [A=71] $\mathbf{c=4.559,\,t=2.3}$ 0.0681
Table 1: Comparison of different electron densities for the $1s$ bound state of 71Ge at a specific radius $r_{0}$, $4\pi\,|\psi^{\rm b}_{e,1s}(r_{0})|^{2}$, according to various calculations available in the literature, which consider nuclear charge densities, $\rho_{\rm ch}(r)$. “A” refers to the considered germanium mass number, which typically enters in the definition of the charge radius. The table presents results from various references, specifying $\rho_{\rm ch}(r)$ (box or 2pF distribution), the associated input parameters, and the corresponding values of $4\pi\,|\psi^{\rm b}_{e,1s}(r_{0})|^{2}$. We report the electron density as obtained using a different Dirac equation solver, GRASP [55]. In the last seven rows, we show the values obtained in this work for $r_{0}=0$, $r_{0}=\sqrt{5/3}R_{\rm ch}$ and with the averaging procedure (avg) considering different values of $C_{\rm ex}$ entering Eq. 13.
DHFS calculations neglect additional effects such as the Breit interaction, vacuum polarization and self-energy corrections, which however are expected to be subdominant for the problem under investigation. In Fig. 1, we compare our results based on RADIAL to those obtained exploiting the GRASP [55] software as well as to other results available in the literature in terms of $\rho_{\rm 1s}(r)=2|\psi^{\rm b}_{e,1s}(r)|^{2}$, where the factor two represents the electron multiplicity. GRASP solves the Dirac equation without any approximation for the exchange effects, including also the aforementioned contributions. Interestingly, the results obtained using RADIAL for $C_{\rm ex}=1$ are in perfect agreement with those of GRASP. Moreover, we observe that a variation of $C_{\rm ex}$ produces a difference in the electron density at the nucleus smaller than 1%, in agreement with the findings of Ref. [57]. Given that $C_{\rm ex}$ is expected to assume a value between $C_{\text{ex}}=3/2$ and $C_{\text{ex}}=1$, we chose to fix its values to the average one, namely $C_{\rm ex}=1.25$, adding a systematic uncertainty to account for the observed spread. The density at the nucleus corresponding to this choice is also shown in Fig. 1, from which it can be compared to the Dirac-Fock result obtained in Refs. [58, 57], finding an excellent agreement. The numerical values obtained for the electron density are reported in Tab. 1 in units of $(\hbar/m_{e}c)^{3}$, along with those available in the literature. We checked that, when using similar assumptions, the electron densities evaluated with our code are compatible with those reported in the literature within about 0.2%.
Besides the numerical evaluation of the electron density at the nucleus, we employed the same dedicated software, based on the RADIAL package [53], also to numerically evaluate the free electron wave functions for a finite-size nucleus, which enter the generalised definition of the Fermi function in Eq. (10). This has to be compared to the common definition of the Fermi function which can be found in the literature (see e.g. Ref. [27]), obtained by the product of various contributions, namely
@@
with $\gamma\equiv\sqrt{1-(\alpha Z)^{2}}$ and $\eta\equiv(\alpha ZE_{e})/p_{e}$, where $\alpha$ is the fine-structure constant. The additional term $L_{0}(E_{e},Z)$ corrects for the finite size of the nucleus considering a box density with equivalent radius $R_{\rm box}=\sqrt{5/3}R_{\rm ch}$, while $U(E_{e},Z)$ is introduced to improve the nuclear distribution from a box density to a more realistic Fermi distribution. Finally, $S(E_{e},Z)$ introduces a correction to the outgoing wave function due to atomic screening [54, 65, 53]. We have checked that by considering a nucleus with a uniform density and evaluating the Fermi function at the origin, we retrieve the tabulated values of the Fermi function from Behrens et al. [38], expressed as the first two terms of Eq. (14), namely $F_{0}L_{0}$.
Figure 2: Fermi function for 71Ge as a function of the radial distance (blue solid line), as derived in this work in Eq. (10) using the exact DHFS $g_{\kappa}(r)$ and $f_{\kappa}(r)$ wave functions for $E_{e}=1.031\,\mathrm{MeV}$. We also show the charge density of 71Ge, $\rho_{\rm ch}(r)$ (gold dashed line) in arbitrary units and the electron density for a bound electron in the 1ss state of $^{71}\text{Ge}$ (red long dashed line), as obtained from our numerical code for $C_{\rm ex}=1.25$. The vertical dotted and dash-dotted lines represent $R_{\mathrm{ch}}$ and $R_{\mathrm{box}}=\sqrt{5/3}\,R_{\rm ch}$, respectively.
@@
In Fig. 2, we show our result for the Fermi function calculated using the exact DHFS $g_{\kappa}(r)$ and $f_{\kappa}(r)$ wave functions in Eq. (10) for $E_{e}=1.031\,\mathrm{MeV}$. Our result has been obtained by calculating all of the contributions to the Fermi function within a unique and coherent framework and is in agreement within 0.2-0.3% with the values reported in Ref. [27], as shown in Table 2 for $r_{0}=0$. Moreover, we have checked that the choice of $C_{\rm ex}$ does not affect the free electron wave functions, with an impact on the evaluation of the Fermi function below 0.004%. In Fig. 2, we also show the electron density for the $1s$ bound state of 71Ge as obtained by our numerical calculation for $C_{\rm ex}=1.25$ and the 2pF distribution used to describe the nuclear density.
@@
As a further improvement, we move beyond the conventional approximation of evaluating the electron wave function at a single point, as done in Eqs. (2) and (5). We propose the adoption of averaged wave functions over the $^{71}{\textrm{Ge}}$ nuclear charge density, in both the calculations of IBD and EC amplitudes. [^7] The average should be evaluated over the Gamow-Teller transition density, defined as the overlap between the 71Ge and 71Ga nuclear wave functions with the interaction Hamiltonian. However, this quantity is not currently accessible from first principles in nuclear theory, and we leave its inclusion for an upcoming work. This modification also results in an averaging procedure of the Fermi function in Eq. (10), $\overline{\mathcal{F}}(E^{j}_{e},Z_{\text{Ge}})$, consistent with the prescription originally suggested by Bahcall [35] for a point-like nucleus. [^8] Bahcall suggested to apply the averaging procedure to $F_{0}$ considering that the “electron-capture can occur anywhere within the nuclear volume” [35]. This is not completely justified, since in the evaluation of $F_{0}$ the nucleus is assumed to be point-like and considering that the Fermi function corrects the wave function of the outgoing electron and it does not refer to the electron captured in the EC process. On the contrary, the introduction of exact DHFS electron wave functions, as done in this work, naturally leads to the evaluation of the average of the latter within the nucleus when factoring out the electron wave function from the matrix element as in Eq. (2). The averaging is done through
$\overline{\mathcal{F}}(E_{e},Z)=\frac{\int{\rho_{\rm ch}(r)\,\mathcal{F}}(E_{e},Z,r)\,\mathrm{d}^{3}r}{\int{\rho_{\rm ch}(r)\,\mathrm{d}^{3}r}}\,$, (16)
@@
The effect of averaging the electron density is shown in the last row of Table 1, where we compare the values obtained at a fixed radius, namely $r_{0}=0$ and $r_{0}=\sqrt{5/3}R_{\rm ch}$, with that retrieved after the averaging procedure.
@@
51Cr 1.031 2.774 2.782 2.732
1.030 2.775 2.782 2.733
1.025 2.779 2.786 2.736
0.711 3.306 3.314 3.257
0.710 3.309 3.317 3.260
0.705 3.330 3.338 3.281
37Ar 1.092 2.734 2.741 2.692
1.092 2.734 2.742 2.692
1.089 2.736 2.743 2.294
Table 2: Comparison between the Fermi function defined according to Eq. (14[27], $F(E_{e},Z_{\text{Ge}})$, and those calculated with our software according to Eqs. (10) and (16), $\mathcal{F}(E_{e},Z_{\text{Ge}})$ and $\overline{\mathcal{F}}(E_{e},Z_{\text{Ge}})$, for electron energies $E_{e}$ relevant to $^{51}\text{Cr}$ and $^{37}\text{Ar}$ neutrino sources. In the second and third columns we use $r_{0}=0$.
@@
To calculate $f_{\text{EC}}$, following Eq. (7), we keep into account the so-called overlap and exchange correction, $\epsilon_{o}^{1s}$, given that we write the total rate in terms of the 1s-capture rate. We compute the average of the values reported in Refs. [68, 69], as previously done in Ref. [27], yielding to $\epsilon_{o}^{1s}=-0.013\pm 0.007$. We also consider an additional term due to the weak magnetism correction, $1+\epsilon_{\text{q}}]_{\text{EC}}$, with $\epsilon_{\text{q}}=(4.9\pm 0.7)\,\times 10^{-4}$ [27].
@@
$B\mathrm{{}_{GT}}(\mathrm{gs})=\frac{|\mathcal{M}_{\rm nuc}^{\rm{IBD}}|^{2}}{2J_{\rm Ga}+1}=0.0859\pm 0.0006\,$, (23)
@@
In Table 3, we report the ground-state cross section for the IBD process for 51Cr and 37Ar neutrino sources. The average procedure reduces the ground state cross section by 2.8%, indicating a non-negligible correction that should be taken into account. Moreover, the reported uncertainty on the cross section includes also a systematic uncertainty due to the choice of the exchange potential parameter, estimated as the maximum difference between the cross section calculated for $C_{\rm ex}=1$ and $C_{\rm ex}=3/2$ with respect to that obtained for our benchmark value of $C_{\rm ex}=1.25$. This contribution corresponds to about 0.4%.